OpenStudy (anonymous):

sqrt0.5 - 2sqrt1/3 - sqrt1/8 + sqrt75

OpenStudy (anonymous):

\[\sqrt{0.5} - 2\sqrt{1/3} - \sqrt{1/8} + \sqrt{75}\]

OpenStudy (anonymous):

= ?